OpenStudy (anonymous):

Hello. i am Studying for a math test help will be rewarded i dont understand how to do this one equation.

OpenStudy (anonymous):

what equation

OpenStudy (anonymous):

wrong one

OpenStudy (anonymous):

ok

OpenStudy (kaylardz0405):

wouldnt you have to make a dot plot

OpenStudy (kaylardz0405):

and see wich way the the information is skwed

OpenStudy (anonymous):

i think

OpenStudy (unknownunknown):

Total narrowed notebooks = 102+67+40= 209 Those 1-subject narrow ruled = 102 102/209 = 48.8%

OpenStudy (anonymous):

Thank you !!

OpenStudy (unknownunknown):

Np

OpenStudy (anonymous):

0.4880382775119617 but if you 0.488=48%8